OpenStudy (anonymous):

I need help expanding this binomial (2x^2 + y^2)^4

OpenStudy (welshfella):

use the identity (a + b)^n = a^n + n a^(n-1) b + n(n-1) a^(n-2) b^2 +........ + b^n ----- 2!

OpenStudy (welshfella):

I'll start if off for you a = 2x^2, b = y^2 and n = 4 (2x^2 + y^2)^4 = (2x^2)^4 + 4* (2x^2)^3* y^2 +